OpenStudy (anonymous):

One number is 50 more than twice another. Their product is 50 more than twice their sum. Find the numbers

OpenStudy (triciaal):

let the numbers be x and y y = 2x + 50 x*y = 50 + 2(x + y) more than means add twice means 2 times product means multiply and sum means add you have 2 equations you need to solve simultaneously hint 2x = 2x solve for x substitute in an original to find y

OpenStudy (anonymous):

im trying to work this

OpenStudy (triciaal):

I did it twice and got x = 3 so that means y = 56

OpenStudy (triciaal):

what do you have?

OpenStudy (anonymous):

I have to come up with four numbers. A set of lower values and higher values set

OpenStudy (triciaal):

I am thinking that I used x*y also need to use (-x)*(-y) absolute value somewhere

OpenStudy (triciaal):

I am thinking use when x > y need to consider x< y

OpenStudy (triciaal):

what if any values do you have?

OpenStudy (triciaal):

@mathstudent55 do you mind helping?

Directrix (directrix):

y = 2x + 50 x*y = 50 + 2(x + y) --------------------------- x(2x + 50) = 50 + 2(x + (2x + 50) ) 2x^2 + 50 x = 50 + 2( 3x +50) 2x^2 + 50 x = 50 + 6x + 100 2x^2 + 50 x = 6x + 150 2x^2 + 50 x - 6x = 150 2x^2 + 44x - 150 = 0 x^2 +22x - 75 = 0 ( x + 25 ) (x - 3) = 0 **** x = - 25 OR x = 3 @chris @triciaal

Directrix (directrix):

x = - 25 OR x = 3 If x = - 25, then y = 2*(-25) + 50 = 0 Result One: ( - 25, 0 ) ----> If x = 3, then y = 2*(3) + 50 = 56 Result Two: (3, 56)

OpenStudy (triciaal):

@Directrix thanks when y = 0 then 0-2x= 50 so x = -25 by the way equating the (2x) or the (50) was much easier and faster than substitution for y

Directrix (directrix):

"Fast is fine but accuracy is everything." ~Wyatt Earp The reason I attempted to help in this thread was to assist in finding the missing two answers, the absence of which was mentioned above. Result One: ( - 25, 0 ) ----> If x = 3, then y = 2*(3) + 50 = 56 Result Two: (3, 56)